OpenStudy (goformit100):

Choose the last step in the given solution to the inequality 5x - 14 greater than or equal to 7x + 10. 5x - 14 greater than or equal to 7x + 10 -2x - 14 greater than or equal to 10 -2x greater than or equal to 24 x greater than or equal to -12 x greater than or equal to 12 x less than or greater to 12 x less than or greater to -12

OpenStudy (goformit100):

@ivettef365

OpenStudy (goformit100):

@Fifciol

OpenStudy (goformit100):

@Luis_Rivera

OpenStudy (anonymous):

5x-14>=7x+10 -2x-14>=10 -2x>=24 x<=-12

OpenStudy (ivettef365):

the next step is to divide by -2, since you are dividing by a negative the sign switch

OpenStudy (ivettef365):

don't see any choice that makes sense, should be x less than or equal to -12